a set of forces are acting on an object. The first force, with a magnitude of 24.70N, is directed at -258 degrees and the second force, with a magnitude of 53.20B, is directed at 37 degrees north of east. The third force with a magnitude of 166.50N is directed at 270, and the fourth force of 49.92N is directed at 50 degrees east of south. Solve the resultant force and exact direction analytically.


1
Expert's answer
2021-02-04T17:13:26-0500

"F_x=-F_1cos78\u00b0+F_2cos37\u00b0-F_3+F_4sin50\u00b0=-80.64~N."

"F_y=F_1sin78\u00b0+F_2sin37\u00b0-F_4cos50\u00b0=24.09~N."

"F=\sqrt{F_x^2+F_y^2}=84.16~N" directed at "arctan\frac{F_x}{F_y}=73.37°" west of north.